# Quantum Monte Carlo simulation of light and superlight bipolarons in extended Hubbard–Holstein models on face- and body-centered-cubic lattices

### v1

@@ 3 Results
Figure 3: Phase diagram of the extended Holstein model on BCC and FCC lattices shown for three different values of $\gamma=1-\Phi_{\rm NN}/\Phi_{00}$. The additional intersite interaction promotes S1 pairing leading to an increase in the size of the S1 region such that it can be found at lower $U$ and $\lambda$. The case $\gamma=1$ corresponds to the Holstein interaction. Lines are a guide to the eye, determined by fitting a cubic curve to the points defining the S0 boundary.
The phase (binding) diagram of the Holstein and extended Holstein bipolarons is shown in Fig. 3. The S0 part of the phase diagram was determined by searching for a sudden drop in the number of phonons associated with the two particles as $U$ is increased at fixed $\lambda$ followed by a region where there is no change in $N_{\rm ph}$ on further increase in $U$. For the regions of parameter space with no S0 bipolaron, the S1 bipolaron is identified by determining where the particle separation is the intersite spacing, $b$ and unbound cases where the particle separation is very large. For FCC lattice, $b=a/\sqrt{2}$ and for BCC lattice, $b=a\sqrt{3}/2$. The Holstein cases are shown in panels (a) and (d). We have not identified an S1 region in this case, although we note that a tiny S1 region can be observed for the 2D Holstein bipolaron [17].
Figure 4: Number of phonons associated with the bipolaron. An abrupt change in the number of phonons corresponds to the boundary between S0 and S1 bipolarons. There are approximately 50% more phonons associated with the FCC lattice due to the larger kinetic energy relative to $\hbar\omega$ in that case. $N_{\rm ph}$ is constant on change of $U$ for S1 and unbound cases.
@@
Figure 5: Inverse radius of the bipolaron, $a/R$. The S1 bipolaron is characterized by pairs with interparticle spacing, $b$ at large $U$ and large $\lambda$. The S0 bipolaron is small and characterized by large inverse radius (found for large $\lambda$ and small $U$). Regions with no pairing are found for large $U$ and small $\lambda$ and shown in black.
@@
Figure 6: Inverse mass of the bipolaron in units of the electron mass, $m_{0}/m^{**}$. $\hbar\omega=t$ in both cases. A sharp change is found at the boundary of the S0 region of the binding diagram.
@@
Figure 7: Transition temperature estimates. The FCC bipolaron has a higher maximum transition temperature than the BCC case for $\gamma=0.5$, in spite of the larger polaron mass. For $\gamma=0.75$, the BCC case has a higher transition temperature. There is a small region of higher $T_{C}$ associated with large bipolarons at low $U$ and $\lambda$ for $\gamma=1$ (Holstein model) that is outside the BEC regime.
@@
Figure 8: Inverse mass, $N_{\rm ph}$, inverse radius and estimated transition temperature of the FCC and BCC bipolarons at large $\lambda$ and $\hbar\omega=W$. The FCC bipolaron is lighter (superlight) in this limit, with the S1 bipolaron much lighter than the S0 bipolaron in all cases. The small peak in inverse mass for $\lambda=2$ and BCC lattice (and to a lesser extent for the FCC lattice) corresponds to the light bipolaron case where intra and inter-site effective interactions are of similar size on the boundary between S0 and S1 bipolarons. For larger phonon frequency, the number of phonons changes more slowly at the boundary between S0 and S1 bipolarons. Transition temperature increases monotonically with $U$. $\gamma=0.5$. As $\lambda$ increases, the percentage difference between the BCC and FCC cases becomes larger as the superlight regime is approached. For each data point, the error bars are displayed as 3 standard deviations. Where the error bars are not visible it means that they are too small to be apparent on the scale of the plot.
@@
Figure 8 shows the inverse mass, number of phonons, inverse radius and estimated transition temperature of the FCC and BCC bipolarons at large $\lambda$ and $\hbar\omega=W$. The FCC bipolaron is lighter than the BCC case by around 50% in this limit, with the S1 bipolaron much lighter than the S0 bipolaron in all cases. At the larger $\lambda$, the bipolaron is slightly lighter at the boundary between the S0 and S1 bipolarons, as a manifestation of the crawler resonance of the bipolaron which can be identified in the antiadiabatic limit (where the on-site and intersite bipolaron configurations have similar energy) as in Fig. 1(B). The superlight mechanisms that we have previously discussed have been established for the anti-adiabatic limit of high phonon frequency. We therefore expect systems with large phonon frequencies (typically the case in systems with low kinetic energies) to have more signatures of this superlight mechanism. At phonon frequency $\hbar\omega=W$, the number of phonons associated with the bipolaron shows a smoother crossover between the S0 and S1 states. The number of phonons is much smaller than for the $\hbar\omega=t$ case reflecting the additional energy carried with the phonons in the bipolaron cloud. The inverse radius displays similar behaviour to the $\hbar\omega=t$ case.
@@
Finally, the estimated transition temperature is shown in Panel 8(d). The estimated transition temperature is much higher in the FCC case than the BCC case, and the ratio of the transition temperatures for the two cases increases with the electron-phonon coupling strength as the bipolaron becomes more tightly bound between sites and the superlight effect becomes more important. The bipolarons become lighter at the point where the on-site and intersite effective interactions become similar. We do not notice a peak in the estimated transition temperature associated with $U^{\prime}\sim V^{\prime}$ in either the BCC or FCC cases.
@@ 4 Discussion and conclusions
Using continuous time path-integral QMC simulations, we have studied the bipolaron in the Hubbard–Holstein and extended Hubbard–Holstein models, which contain local Coulomb repulsion in the presence of phonon-mediated attractive interactions. We have focussed on pairing in BCC and FCC lattices. We used $N_{\rm ph}$ and $R$ to determine the binding (phase) diagram. Mass was calculated using twisted boundary conditions. Transition temperatures were estimated. A benefit of the QMC simulation is that we can examine the effect of phonon retardation on the properties of bipolarons. When $\hbar\omega\ll W$, the retardation effects are significant.
@@
Overall, we found no qualitative difference in bipolaron binding diagrams for BCC and FCC lattices for $\hbar\omega=t$. S1 bipolarons at intermediate electron-phonon coupling tend to be lighter on the FCC lattice for larger intersite coupling and heavier for lower intersite coupling. In the Holstein case, the FCC bipolaron is much heavier and as $\gamma$ decreases, the FCC is lighter.
@@
On increase of phonon frequency, retardation effects become less important and the effective interaction Hamiltonian becomes more Hubbard like. We observed that for $\hbar\omega=W$, signatures of superlight states (where the pair moves by first-order hops) start to emerge, and lower masses are found for FCC lattices, especially at large $\lambda$. We also see evidence of superlight behaviour for pairs on BCC lattices of the type shown in Fig. 1(B) for the specific case where $U^{\prime}=V^{\prime}$. Estimated transition temperatures in this case are larger for bipolarons on FCC lattices.
@@
As further work, we think study of the EPIs in A3C60 compounds would be interesting to establish whether any longer range EPIs that could lead to superlight effects are present. Additional QMC calculations including next-nearest neighbour hopping would also be of interest for study of the BCC lattice, since the intersite spacings of near-neighbour and next-nearest neighbour are similar, potentially leading to similar hoppings with other large effects [15]. This is beyond the scope of the current paper as it would require significant algorithmic development with additional 3 kink updates including next-nearest neighbour hopping and is reserved for a future publication.
Appendix A Quantum Monte Carlo: Update Rules and Weighting Scheme
The electron paths are continuous in time and an electron hops between sites via kinks in its path. In an exchange configuration, our result is free from sign problem since we only studied singlet states. The selection of path, insertion or removal of kink, and choosing the kink type is implemented probabilistically. The kink type specifies the direction of particle’s hop.
Kink insertion or removal is carried out in pairs so that the end configurations of the paths maintain their boundary conditions. We refer to kink insertions and/or removals as updates and there are rules for carrying out these updates for the simulation of two particles. Four binary updates were used:
  1. 1.
    Kinks of type $l$ are added to or removed from each path.
  2. 2.
    Add or remove a kink of type $l$ and its anti-kink $-l$ to or from one of the paths.
  3. 3.
    Insert a kink of type $l$ to one path and remove another kink of the same type $l$ from the same path. This update only works if there is at least one kink of type $l$ on that path since there will no kink to remove if it is non-existent.
  4. 4.
    Insert a kink of type $l$ from one path and remove an anti-kink $-l$ from the other path.
All the updates above were the existing rules prior to the commencement of this study. The detailed balance equations, electron-phonon action, and Metropolis-Rosenbluth conditions for these updates can be found in Ref. [16]. A new update rule for FCC lattices is detailed in the main text.
The summary of the probabilities for kink selection is given below.
  • For $N_{k}$ number of nearest-neighbours, we choose a kink type $l$ with the probability $P_{l}=1/N_{k}$ and its anti-kink $l$ is determined.
  • Since there are only two particles, we select one of the two paths as path $A$ with a probability $1/2$ and label the other as path $B$.
  • We choose a kink shift type with an equal probability $P_{\rm Shift}=1/2$. The kink shift can either be top shift (in the direction of $l$) or bottom shift (in the direction of $-l$).
  • To remove the first kink of type $l$ from a path $A$ at imaginary time $\tau$, do so with the probability $1/N_{Al}(I)$, where $I$ is the initial configuration before removal.
  • To remove the second kink of type $l$ from path $A$, this is done with a probability $1/N_{Al}(I)$, where $I$ represents the configuration before removal.
  • The first kink of type $l$ is inserted at time $\tau$ with probability density $p(\tau)=1/\beta$.
  • For the addition of the second kink of type $l$ at time $\tau^{\prime}$, do so with a probability density $p(\tau)=1/\beta$.
  • The shift type of the second kink is dependent on whether it is correlated or anti-correlated with the shift type of the first kink. The inter-path distance is unchanged for correlated insertion but changes for anti-correlated insertion.

### v2

@@ 3 Results
Figure 4: Phase diagram of the extended Holstein model on BCC and FCC lattices shown for three different values of $\gamma=1-\Phi_{\rm NN}/\Phi_{00}$. Calculations are in the adiabatic limit, $\hbar\omega=t$, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases. The additional intersite interaction promotes S1 pairing leading to an increase in the size of the S1 region such that it can be found at lower $U$ and $\lambda$. The case $\gamma=1$ corresponds to the Holstein interaction. Lines are a guide to the eye, determined by fitting a cubic curve to the points defining the S0 boundary.
3.1 Adiabatic limit
In this section, the adiabatic limit of, $\hbar\omega=t$ is explored, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases.
The phase (binding) diagram of the Holstein and extended Holstein bipolarons is shown in Fig. 4. The S0 part of the phase diagram was determined by searching for a sudden drop in the number of phonons associated with the two particles as $U$ is increased at fixed $\lambda$ followed by a region where there is no change in $N_{\rm ph}$ on further increase in $U$. For the regions of parameter space with no S0 bipolaron, the S1 bipolaron is identified by determining where the particle separation is the intersite spacing, $b$ and unbound cases where the particle separation is very large. For FCC lattice, $b=a/\sqrt{2}$ and for BCC lattice, $b=a\sqrt{3}/2$. The Holstein cases are shown in panels (a) and (d). We have not identified an S1 region in this case, although we note that a tiny S1 region can be observed for the 2D Holstein bipolaron [macridin2004].
Figure 5: Number of phonons associated with the bipolaron. Calculations are in the adiabatic limit, $\hbar\omega=t$, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases. An abrupt change in the number of phonons corresponds to the boundary between S0 and S1 bipolarons. There are approximately 50% more phonons associated with the FCC lattice due to the larger kinetic energy relative to $\hbar\omega$ in that case. $N_{\rm ph}$ is constant on change of $U$ for S1 and unbound cases.
@@ 3.1 Adiabatic limit
Figure 6: Inverse radius of the bipolaron, $a/R$. The S1 bipolaron is characterized by pairs with interparticle spacing, $b$ at large $U$ and large $\lambda$. Calculations are in the adiabatic limit, $\hbar\omega=t$, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases. The S0 bipolaron is small and characterized by large inverse radius (found for large $\lambda$ and small $U$). Regions with no pairing are found for large $U$ and small $\lambda$ and shown in black.
@@
Figure 7: Inverse mass of the bipolaron in units of the electron mass, $m_{0}/m^{**}$. Calculations are in the adiabatic limit, $\hbar\omega=t$, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases. A sharp change is found at the boundary of the S0 region of the binding diagram.
@@
Figure 8: Close packing transition temperature ($T^{*}$) estimates in the adiabatic limit (from Eq. 6) $\hbar\omega=t$, which corresponds to $\hbar\omega=W/12$ for BCC and $W/16$ for FCC cases. The FCC bipolaron has a higher maximum transition temperature than the BCC case for $\gamma=0.5$, in spite of the larger polaron mass. For $\gamma=0.75$, the BCC case has a higher transition temperature. There is a small region of higher $T^{*}$ associated with large bipolarons at low $U$ and $\lambda$ for $\gamma=1$ (Holstein model) that is outside the BEC regime.
@@
Figure 9: Inverse mass of the FCC and BCC bipolarons at large $\lambda$ and intermediate phonon frequency $\hbar\omega=W$. The FCC bipolaron is lighter (superlight) in this limit, with the S1 bipolaron much lighter than the S0 bipolaron in all cases. The small peak in inverse mass for $\lambda=2$ and BCC lattice (and to a lesser extent for the FCC lattice) corresponds to the light bipolaron case where intra- and inter-site effective interactions are of similar size on the boundary between S0 and S1 bipolarons. For larger phonon frequency, the number of phonons changes more slowly at the boundary between S0 and S1 bipolarons. The close packing transition temperature (from Eq. 6) increases monotonically with $U$. As $\lambda$ increases, the percentage difference between the BCC and FCC cases becomes larger as the superlight regime is approached. For comparison, we also include data for the simple cubic (SC) lattice at $\lambda=2$ (see Ref. [davenport2012]), for which all properties are very similar to the BCC case, $a/R$ differs primarily due to the difference in the near-neighbour distance, as does $T^{*}$. Overall, $T^{*}$ is much higher for the FCC case. For each data point, the error bars are displayed as 3 standard deviations. Where the error bars are not visible it means that they are too small to be apparent on the scale of the plot. Lines are a guide to the eye.
3.2 Intermediate adiabaticity
@@
Figure 9 shows the inverse mass, number of phonons, inverse radius and estimated transition temperature of the FCC and BCC bipolarons at large $\lambda$ and intermediate adiabaticity, $\hbar\omega=W$, for $\gamma=0.5$. The FCC bipolaron is lighter than the BCC case by around 50% in this limit, with the S1 bipolaron much lighter than the S0 bipolaron in all cases. At the larger $\lambda$, the bipolaron is slightly lighter at the boundary between the S0 and S1 bipolarons (between small and large $U$), as a manifestation of the crawler resonance of the bipolaron which can be identified in the anti-adiabatic limit (where the on-site and intersite bipolaron configurations have similar energy) as in Fig. 1(B). The superlight mechanisms that we have previously discussed have been established for the anti-adiabatic limit of high phonon frequency. We therefore expect systems with large phonon frequencies (typically the case in systems with low kinetic energies) to have more signatures of this superlight mechanism. At phonon frequency $\hbar\omega=W$, the number of phonons associated with the bipolaron shows a smoother crossover between the S0 and S1 states. The number of phonons is much smaller than for the $\hbar\omega=t$ case reflecting the additional energy carried with the phonons in the bipolaron cloud. The inverse radius displays similar behaviour to the $\hbar\omega=t$ case. For comparison, we also include data for the simple cubic (SC) lattice at $\lambda=2$ (see Ref. [davenport2012]), for which $N_{\rm ph}$ and $m_{0}/m^{**}$ are very similar. $a/R$ differs primarily due to the difference in the near-neighbour distance, as does $T^{*}$. Overall, $T^{*}$ is much higher for the FCC case.
@@
Finally, the close packing transition temperature is shown in Panel 9(d). The estimated close packing transition temperature (from Eq. (6)) is much higher in the FCC case than the BCC case (and SC case), and the ratio of the transition temperatures for the two cases increases with the electron-phonon coupling strength as the bipolaron becomes more tightly bound between sites and the superlight effect becomes more important. The bipolarons become lighter at the point where the on-site and intersite effective interactions become similar, corresponding to a peak in the estimated transition temperature associated with $U^{\prime}\sim V^{\prime}$ in all cases.
@@ 4 Discussion and conclusions
Using continuous time path-integral QMC simulations, we have studied the bipolaron in the Hubbard–Holstein and extended Hubbard–Holstein models, which contain local Coulomb repulsion in the presence of phonon-mediated attractive interactions. We have focused on pairing in BCC and FCC lattices. We used $N_{\rm ph}$ and $R$ to determine the binding (phase) diagram. Mass was calculated using twisted boundary conditions. Transition temperatures were estimated. A benefit of the QMC simulation is that we can examine the effect of phonon retardation on the properties of bipolarons. When $\hbar\omega\ll W$, the retardation effects are significant.
@@
Overall, we found no qualitative difference in bipolaron binding diagrams for BCC and FCC lattices for $\hbar\omega=t$. S1 bipolarons at intermediate electron-phonon coupling tend to be lighter on the FCC lattice for larger intersite coupling and heavier for lower intersite coupling. In the Holstein case, the FCC bipolaron is much heavier and as $\gamma$ decreases, the FCC is lighter. We note that there are also fulleride systems with simple cubic (SC) lattices. We have previously studied the SC version of the extended Holstein Hamiltonian [davenport2012] for $\hbar\omega=W$.
@@
On increase of phonon frequency, retardation effects become less important and the effective interaction Hamiltonian becomes more Hubbard like. We observed that for $\hbar\omega=W$, signatures of superlight states (where the pair moves by first-order hops) start to emerge, and lower masses are found for FCC lattices, especially at large $\lambda$. We also see evidence of superlight behaviour for pairs on BCC lattices of the type shown in Fig. 2(B) for the specific case where $U^{\prime}\sim V^{\prime}$. Estimated transition temperatures in this case are larger for bipolarons on FCC lattices.
As further work, study of the EPIs in A3C60 compounds would be interesting to establish whether any longer range EPIs that could lead to superlight effects are present in fullerides. We briefly discuss our results in the context of parameters relating to the A3C60 compounds. These materials display superconductivity for both BCC and FCC lattice structures. Their bandwidth $W\sim 0.5\mathrm{eV}$. The Coulomb repulsion in A3C60 superconductors is also large with $1.5\lesssim U/W\lesssim 2.5$ and typical values for the nearest-neighbour Coulomb potentials range from about 0.50.8W0.5-0.8W [gunnarsson1996mott, gunnarsson2004alkali, lof1992band, martin_and_ritchie_1993, antropov1992coulomb, lof1992correlation, bruhwiler1993auger]. The high-frequency intramolecular vibron modes, $\hbar\omega/W\sim 0.1-0.4$, play a key role in the high critical temperatures. Electron-phonon coupling involves multiple bands (of a Jahn–Teller type), but single band models may be used to explore some of the physics involved with the electron-phonon interactions. The strongest electron-phonon couplings are intra-molecular and involve coupling between electrons and Einstein modes. They are estimated from experiments to be of order $\lambda\sim 0.5-1.0$. The different types of electron-phonon interaction present in the fullerides are described in Ref. [gunnarsson2004alkali]. Intersite electron-phonon interactions mediated by neighbouring $C_{60}$ sites or alkali metal sites would be required to obtain S1 bipolarons. Inter-molecular interactions between electrons on C60 molecules are weak. The strength of interactions between electrons on C60 molecules and alkali metal ions is unclear (with DFT calculations suggesting $\lambda\lesssim 0.1$, but other calculations estimating higher interactions). Determining the strength of these interaction in more detail would be useful to establish if the they are sufficiently strong to bind lighter S1 bipolarons.
@@
Additional QMC calculations including next-nearest neighbour hopping would also be of interest for study of the BCC lattice, since the intersite spacings of near-neighbour and next-nearest neighbour are similar, potentially leading to similar hoppings. The presence of similar hopping strengths can lead to superlight effects in the anti-adiabatic limit, and the results in this paper indicate that superlight effects can persist into the intermediate adiabicity regimes [adebanjo2024]. In practice, these next-nearest neighbor interactions are likely to decrease mass and increase transition temperatures due to the 1st order processes for bipolaron motion enabled by those extra hops. We discussed the process in the anti-adiabatic limit in our recent paper [adebanjo2024]. QMC simulation of the effects of the next-nearest neighbour hoppings is beyond the scope of the current paper as it would require significant algorithmic development with additional 3 kink updates involving both nearest and next-nearest neighbour hopping and is reserved for a future publication.